# M-LRM: Multi-view Large Reconstruction Model

### v1

@@ Abstract
Despite recent advancements in the Large Reconstruction Model (LRM) demonstrating impressive results, when extending its input from single image to multiple images, it exhibits inefficiencies, subpar geometric and texture quality, as well as slower convergence speed than expected. It is attributed to that, LRM formulates 3D reconstruction as a naive images-to-3D translation problem, ignoring the strong 3D coherence among the input images. In this paper, we propose a Multi-view Large Reconstruction Model (M-LRM) designed to efficiently reconstruct high-quality 3D shapes from multi-views in a 3D-aware manner. Specifically, we introduce a multi-view consistent cross-attention scheme to enable M-LRM to accurately query information from the input images. Moreover, we employ the 3D priors of the input multi-view images to initialize the tri-plane tokens. Compared to LRM, the proposed M-LRM can produce a tri-plane NeRF with $128\times 128$ resolution and generate 3D shapes of high fidelity. Experimental studies demonstrate that our model achieves a significant performance gain and faster training convergence than LRM. Project page: https://murphylmf.github.io/M-LRM/
Figure 1: Our M-LRM can generate novel views of high fidelity and meshes of high quality given single or multiple views as input.
footnotetext: Corresponding authors.
1 Introduction
@@
  • Our M-LRM addresses the shortcomings of instant3D, which lacks modeling of spatial correlation among multiple input views. This leads to slow convergence and sub-optimal generation results. Instead, our M-LRM leverages the spatial coherence of input images to efficiently generate high-fidelity results.
  • We propose a novel Geometry-aware Positional Embedding mechanism to fully utilize image prior. This helps the randomly initialized tri-plane capture spatial and image priors, enabling faster convergence of our M-LRM.
@@
  • Through extensive experiments, we demonstrate that our M-LRM outperforms all baseline methods in sparse-view reconstruction tasks. It can also generate high-quality 3D objects with a single image input when equipped with a multi-view generation model like Zero123++ shi2023zero123++ .
2 Related Work
Figure 2: Visualization of intermediate results during training process. Compared with existing LRM-based approaches, like Instant3D li2023instant3d , our proposed GCA and GaPE remarkably reduce the convergence time and improve the reconstruction quality by a large margin.
@@ 2 Related Work
3D Generation. Early methods on image-to-3D mainly employ Generative Adversarial Networks (GANs) to produce 3D representations, like point cloud li2018point , signed distance function (SDF) shi2022deep ; zou2023triplane ; chan2022efficient or neural radiance field (NeRF) gu2021stylenerf ; niemeyer2021giraffe . By incorporating 2D GAN architectures brock2018large ; creswell2018generative ; kang2023scaling ; karras2017progressive ; karras2021alias ; karras2019style ; karras2020analyzing ; zhu2023exploring , 3D GANs have made remarkable advancements to generate 3D-aware content in a feed-forward manner. As a representative attempt, EG3D chan2022efficient employs GANs to generate samples in latent space and introduce a compact triplane representation for rendering novel views of 3D objects. Another research branch in 3D generation utilizes diffusion models (DMs) tewari2024diffusion ; ho2020denoising ; rombach2022high ; song2020score ; po2023state ; ddim ; ddpm ; ldm , which have emerged with unprecedented success. One approach involves training 3D DMs directly using 3D supervision gupta20233dgen ; jun2023shap ; nichol2022point ; ntavelis2023autodecoding ; shue20233d or 2D supervision anciukevivcius2023renderdiffusion ; chen2023single ; gu2023nerfdiff ; karnewar2023holodiffusion ; liu2023zero ; shen2023gina . However, these approaches suffer from 3D inconsistency and poor quality due to the lack of 3D priors. DreamFusion poole2022dreamfusion introduces the Score Distillation Sampling (SDS) loss and generates NeRF under the supervision of pretrained large-scale 2D DMs. Variants of DreamFusion chen2023fantasia3d ; chung2023luciddreamer ; hertz2023delta ; liang2023luciddreamer ; lin2023magic3d ; poole2022dreamfusion ; shi2023mvdream ; tang2023dreamgaussian ; wang2023score ; wang2024prolificdreamer aim to reduce Janus (multi-faced) artifacts and improve geometry quality. However, the prolonged optimization process limits its practical application.
@@
Sparse-view Reconstruction Multiview images provide stereo cues facilitating more accurate and detailed shape generation. Consequently, sparse view-based neural reconstruction methods chen2021mvsnerf ; jain2021putting ; lin2023vision ; long2022sparseneus ; wang2021ibrnet ; yu2021pixelnerf achieve significant improvements in generalization to unseen scenes. However, challenges remain in capturing patterns within large-scale datasets. Scaling up the model and datasets is useful, but it also introduces new challenges, such as increased computational complexity. To address these challenges, recent methods leverage large-scale 3D datasets and well-designed multiview diffusion models (DMs) to generate multiple novel views from a single image shi2023mvdream ; wang2023imagedream ; long2023wonder3d ; liu2023syncdreamer ; tang2023MVDiffusion ; tang2024mvdiffusion++ ; liu2023zero123 ; shi2023zero123++ ; one2345 ; li2024era3d . Subsequent steps involve neural reconstruction and Gaussian kerbl20233d or reconstruction tang2024lgm ; xu2024grm . However, these explicit reconstruction methods face challenges due to multi-view inconsistency and low image resolution. Inspired by hong2023lrm ; wang2023pf ; zou2023triplane , we employ the recently proposed transformed-based large reconstruction model for implicit 3D shape generation.
@@
Large Reconstruction Model The extensive 3D datasets deitke2023objaverse ; deitke2024objaverse have opened up new possibilities for training reconstruction models that can create 3D models from 2D images. LRM hong2023lrm first showcased the potential of using the transformer backbone to map image tokens to implicit 3D triplanes with multi-view supervision, resulting in highly generalizable models. Instant3D li2023instant3d achieves highly generalizable and high-quality single-image to 3D generation by extending LRM to a sparse-view reconstruction model and utilizing the prior in 2D diffusion models. While LGM tang2024lgm and GRM xu2024grm models have replaced the triplane NeRF representation with 3D Gaussians kerbl20233d to improve rendering efficiency, Gaussians have limitations in explicit geometry modeling and high-quality surface extraction. Concurrent works $\text{MVD}^{2}$ zheng2024mvd2 , CRM wang2024crm , and InstantMesh xu2024instantmesh use the same transformer backbone and focus on optimizing mesh representation for high-quality geometry and texture modeling. In contrast, our work focuses on incorporating 3D geometry prior to the memory-consuming transformer attention for efficient training. Moreover, our proposed M-LRM, featuring a novel 3D-aware transformer block, allows significant generalization in 3D shape modeling.
3 M-LRM
@@
Figure 3: Overview of M-LRM. M-LRM is a fully differentiable transformer-based framework, featuring a feature encoder, geometry-aware position encoding and a multi-view cross attention block. Given multi-view images with corresponding camera poses, M-LRM incorporates the 2D and 3D features to efficiently conduct 3D-aware multi-view attention. The proposed geometry-aware position encoding allows more detailed and realistic 3D generation.
@@
Fig. 3 illustrates the overall framework of our M-LRM. Unlike LRM which treats single-image-to-3D as a pure image translation problem, our method fully leverages the geometry prior to link 2D image and 3D triplane, thus achieving faster convergence and being able to generate results of better quality. Specifically, we achieve such advances in two steps. (1) Given multiple views of an object, we first extract the image features with an image encoder and propose Geometry-aware Positional Embeddings to initialize the learnable triplane tokens as detailed in Sec. 3.1. (2) With the learnable triplane tokens, we further adopt Geometry-aware Cross Attention to achieve the image feature injection in a geometry-aware manner as described in Sec. 3.2. Finally, we utilize the NeRF mildenhall2021nerf renderer to generate 3D contents with the generated triplane token.
3.1 Geometry-aware Positional Embeddings
@@
In LRM hong2023lrm and its variants wang2023pflrm ; li2023instant3d , the triplane tokens are randomly initialized without any priors and then optimized in the training. For convenience, we denote the randomly initialized triplane tokens as $\mathbf{T}_{r}$. In this section, we propose Geometry-aware Positional Embedding (GaPE) to better utilize image prior instead of the randomly initialized triplane tokens.
@@
Geometry Volume Construction. Inspired by the previous multi-view reconstruction method long2022sparseneus , we aim to directly learn a geometry volume from the input images which encodes local geometric information based on the hints of multi-view consistency of the images. To cope with this, we first construct a grid volume whose vertices $\mathbf{X}=\{\mathbf{x}^{j}\}_{j\in\mathcal{J}}$ are uniformly sampled in the region of interest to be reconstructed. Considering the $i$-th view among the $N$ input views, with its intrinsic $K_{i}\in\mathbb{R}^{3\times 3}$ and extrinsic $E_{i}\in\mathbb{R}^{3\times 4}$, we back project the vertices to the feature map of the $i$-th view $F_{i}$, and derive the corresponding feature grid ${G}_{i}=\{F_{i}(\pi_{i}(\mathbf{x}^{j}))\}_{j\in\mathcal{J}}$ for the $i$-th view, where $\pi_{i}$ indicates the projection that projecting a 3D vertex onto the $i$-th image plane. The projection $\pi_{i}$ is defined as
@@
where $\mathbf{x}^{j}\in\mathbb{R}^{3\times 1}$ is the 3D coordinate of the grid vertex and $\tilde{\mathbf{x}}^{j}\in\mathbb{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{x}^{j}$. $\mathbf{u}^{j}_{i}\in\mathbb{R}^{2\times 1}$ is the 2D coordinate of 2D image space and $\tilde{\mathbf{u}}^{j}_{i}\in\mathbb{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{u}_{i}^{j}$. With the obtained 2D coordinates, we could use $\operatorname{grid\_sample}$ to extract corresponding features from the 2D feature map $F_{i}$. After back-projecting all the grid vertices, we obtain $N$ feature grids $\{G_{i}\}_{i=1,\cdot\cdot\cdot,N}$, which are fused together into one unified feature volume $\mathbf{G}$ with a 3D convolution network.
@@
3.2 Efficient Geometry-aware Cross Attention
Figure 4: The illustration of Geometry-aware Cross Attention.
In addition to the geometry-aware triplane tokens, we introduce a novel cross-attention mechanism, named Geometry-aware Cross Attention (GCA), which incorporates explicit 3D-aware priors for enhanced information fusion between image features and triplane tokens. Specifically, unlike the vanilla attention used in previous LRM works hong2023lrm ; wang2023pflrm ; li2023instant3d that accesses global features from all views, our method constructs an efficient attention mechanism. This mechanism ensures that each triplane token correlates with only a limited number of image feature tokens, improving computational efficiency and relevance.
Our intuition is that each token in a triplane corresponds to a ray in the 3D space passing through the token, which can be projected into each view (as illustrated in Fig. 4). In other words, each token is related to a projected ray within a view rather than the entire view. Consequently, when performing cross-attention between each triplane token and the image features, we only need to correlate the triplane token with the feature tokens intersecting each projected ray. In prior works, dense cross-attention operates on $n\times h\times w$ feature tokens, where $n$ is the number of views, and $h$ and $w$ are the height and width of the feature map. In contrast, our geometry-aware attention operates on only $n\times V$ feature tokens, where $V$ is the size of the geometry volume. This significantly reduces the computational complexity and focuses attention on more relevant features.
@@ 3.2 Efficient Geometry-aware Cross Attention
$Q=W^{Q}t^{s},K=W^{K}\left[f_{e,l}^{s}\right]_{e\in\mathcal{E},l\in\mathcal{L}},V=W^{V}\left[f_{e,l}^{s}\right]_{e\in\mathcal{E},l\in\mathcal{L}},$ (3)
where $\left[f_{e,l}^{s}\right]$ denote the features of the related projected feature tokens, $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable matrices that project the inputs to query, key, and value, respectively.
To preserve global information interaction among different triplane tokens, we only replace half of the original transformer blocks with GCA, as it helps the model to preserve more details.
Figure 5: Illustration of Geometry-aware Cross Attention block (top) and Positional Embeddings (bottom).
3.3 Training Objectives
After we have the corresponding triplane $P$, we adopt differentiable NeRF renderer $\tau_{N}$ to render images and masks as $\{I^{pred},m^{pred}\}=\tau_{N}(E,K|P)$. The predicted novel views are supervised via differentiable rendering by other sampled views. We adopt MSE loss, binary mask loss and LPIPS loss zhang2018unreasonable . The training loss can be formulated as
$\mathcal{L}_{\text{NeRF }}$ $=\lambda_{\text{mse }}\sum_{i}\left\|I_{i}^{\text{pred }}-I_{i}^{\text{gt }}\right\|_{2}+\lambda_{\text{lpips }}\sum_{i}\mathcal{L}_{\text{lpips }}\left(I_{i}^{\text{pred }},I_{i}^{\text{gt }}\right)$ (4)
$+\lambda_{\text{mask }}\sum_{i}\left(m^{gt}\log m^{\text{pred }}+\left(1-m^{gt}\right)\log\left(1-m^{\text{pred }}\right).\right.$
4 Implementation Details and Datasets
Network Architecture. We train two variants of M-LRM, and here we provide details of our base variant. We refer to supplemental materials for more information on tiny variants.
For the image encoder, we utilize DINOv2-base with a patch size of 14. The transformer decoder consists of 12 layers with a hidden dimension of 768. We employ multi-head attention with 8 heads, and each head has a hidden dimension of 64. The learnable triplane tokens are initialized with a resolution of $64\times 64$ and 768 channels. To upsample the triplane, we use a transpose convolution layer with a kernel size of 2 and a stride of 2, resulting in a resolution of 128x128 and 40 channels. The concatenated triplane features are then decoded using a 4-layer MLP into density and RGB values. We sample 192 points per ray.
We construct the spatial index volume and cost volume with a resolution of $64^{3}$. After projecting the image features into spatial volumes, we aggregate them using a 5-layer 3D convolution networkwith channel sizes of [64, 128, 128, 128, 64]. Additionally, we derive additional triplane features from the cost volume using three convolution layers with a kernel size of 3 and a stride of 1.
Other Implementation Details. M-LRM is trained on 32 NVIDIA H800 GPU ( 80GB) for 3 days. The batch size for each GPU was set to 1, and gradient accumulation was employed for 8 steps, resulting in a total batch size of 512. At each iteration, we randomly sample 10 views among the total 24 images, where 6 of them are used as input conditional images, and the other 4 for supervision. AdamW optimizer is used with $\beta_{1},\beta_{2}$ set to $0.9,0.95$ respectively, and weight decay is set to 0.05. We increase the learning rate from 4e-10 to 4e-4 for the first epoch. After warmup, we adopt cosine annealing strategy for the learning rate and the learning rate decay drops to 0 at the end of training. The coefficients for different loss components are $\lambda_{mse}=1.0$, $\lambda_{L1}=0.2$, $\lambda_{mask}=0.1$ and $\lambda_{LPIPS}=1.0$. Following LGM tang2024lgm , we adopt camera augmentation and image grid distortion as data augmentation.

### v2

# Multi-View Large Reconstruction Model via Geometry-Aware Positional Encoding and Attention

@@ Abstract
Despite recent advancements in the Large Reconstruction Model (LRM) demonstrating impressive results, when extending its input from single image to multiple images, it exhibits inefficiencies, subpar geometric and texture quality, as well as slower convergence speed than expected. It is attributed to that, LRM formulates 3D reconstruction as a naive images-to-3D translation problem, ignoring the strong 3D coherence among the input images. In this paper, we propose a Multi-view Large Reconstruction Model (M-LRM) designed to reconstruct high-quality 3D shapes from multi-views in a 3D-aware manner. Specifically, we introduce a multi-view consistent cross-attention scheme to enable M-LRM to accurately query information from the input images. Moreover, we employ the 3D priors of the input multi-view images to initialize the triplane tokens. Compared to previous methods, the proposed M-LRM can generate 3D shapes of high fidelity. Experimental studies demonstrate that our model achieves a significant performance gain and faster training convergence. Project page: https://murphylmf.github.io/M-LRM/.
Index Terms:
3D generation, large reconstruction model, sparse-view reconstruction, single image to 3D.
Figure 1: Our M-LRM is a novel framework that can handle both single-view generation tasks and multi-view (e.g. 6-view) reconstruction tasks. M-LRM facilitates high-quality novel views generation given a single view or multiple views as input.
I Introduction
@@
  • Our M-LRM addresses the shortcomings of instant3D, which lacks modeling of spatial correlation among multiple input views, leading to slow convergence and sub-optimal generation results. Instead, our M-LRM leverages the spatial coherence of input images to generate high-fidelity results.
  • We propose a novel Geometry-aware Positional Embedding mechanism to fully utilize image prior. This helps the randomly initialized triplane capture spatial and image priors, enabling faster convergence of our M-LRM.
@@
  • Extensive experiments demonstrate that our M-LRM outperforms all baseline methods. It can also generate high-quality 3D objects with a single image input when equipped with a multi-view generation model like Zero123++ [1].
II Related Work
II-A 3D Generation
@@ II Related Work
Early methods on image-to-3D mainly employ Generative Adversarial Networks (GANs) to produce 3D representations, like point cloud [2], signed distance function (SDF) [3, 4, 5] or neural radiance field (NeRF) [6, 7]. By incorporating 2D GAN architectures [8, 9, 10, 11, 12, 13, 14, 15], 3D GANs have made remarkable advancements to generate 3D-aware content in a feed-forward manner. As a representative attempt, EG3D [5] employs GANs to generate samples in latent space and introduce a compact triplane representation for rendering novel views of 3D objects. Another research branch in 3D generation utilizes diffusion models (DMs) [16, 17, 18, 19, 20, 21, 22, 23], which have emerged with unprecedented success. One approach involves training 3D DMs directly using 3D supervision [24, 25, 26, 27, 28] or 2D supervision [29, 30, 31, 32, 33, 34]. However, these approaches suffer from 3D inconsistency and poor quality due to the lack of 3D priors. DreamFusion [35] introduces the Score Distillation Sampling (SDS) loss and generates NeRF under the supervision of pretrained large-scale 2D DMs. Variants of DreamFusion [36, 37, 38, 39, 40, 35, 41, 42, 43, 44] aim to reduce Janus (multi-faced) artifacts and improve geometry quality. However, the prolonged optimization process limits its practical application.
II-B Sparse-View Reconstruction
@@
Multi-view images provide stereo cues facilitating more accurate and detailed shape generation. Consequently, sparse view-based neural reconstruction methods [45, 46, 47, 48, 49, 50] achieve significant improvements in generalization to unseen scenes. However, challenges remain in capturing patterns within large-scale datasets. Although scaling up the model and datasets is beneficial, it also introduces new challenges, such as increased computational complexity and sub-optimal generalizability. To address these challenges, recent methods leverage large-scale 3D datasets and well-designed multi-view diffusion models (DMs) to generate multiple novel views from a single image [41, 51, 52, 53, 54, 55, 56, 1, 57]. Subsequent steps involve neural reconstruction or Gaussian reconstruction [58, 59, 60]. However, these explicit reconstruction methods face challenges due to multi-view inconsistency and low image resolution. Inspired by [61, 62, 4], we employ the recently proposed transformed-based large reconstruction model for implicit 3D shape generation.
II-C Large Reconstruction Model
@@
The extensive 3D datasets [63, 64] have opened up new possibilities for training reconstruction models that can create 3D models from 2D images. LRM [61] first showcased the potential of using the transformer backbone to map image tokens to implicit 3D triplanes with multi-view supervision, resulting in highly generalizable models. Instant3D [65] achieves highly generalizable and high-quality single-image to 3D generation by extending LRM to a sparse-view reconstruction model and utilizing the prior in 2D diffusion models. While LGM [59] and GRM [60] models have replaced the triplane NeRF representation with 3D Gaussians [58] to improve rendering efficiency, Gaussians have limitations in explicit geometry modeling and high-quality surface extraction. Concurrent works $\text{MVD}^{2}$ [66], CRM [67], and InstantMesh [68] use the same transformer backbone and focus on optimizing mesh representation for high-quality geometry and texture modeling. In contrast, our work focuses on incorporating 3D geometry prior to the time-consuming transformer attention for efficient training. Moreover, our proposed M-LRM, featuring a novel 3D-aware transformer block, allows significant generalization in 3D shape modeling.
@@
Figure 2: Overview of M-LRM. M-LRM is a fully differentiable transformer-based framework, featuring an image encoder, geometry-aware position encoding and a multi-view cross attention block. Given multi-view images with corresponding camera extrinsic $E$ and intrinsic $K$, M-LRM incorporates the 2D and 3D features to conduct 3D-aware multi-view attention. Our transformer triplane decoder consists of $L$ layers, where each layer is composed of a multi-view cross attention Block and a vanilla transformer block. The proposed geometry-aware position encoding allows more detailed and realistic 3D generation.
III M-LRM
@@
Fig. 2 illustrates the overall framework of M-LRM. Unlike LRM which treats single-image-to-3D as a pure image translation problem, our method fully leverages the geometry prior to link 2D image and 3D triplane, achieving faster convergence and being able to generate results of better quality. Specifically, we achieve such advances in two steps. (1) Given multiple views of an object, we first extract the image features with an image encoder and propose Geometry-aware Positional Embeddings to initialize the learnable triplane tokens. (2) With the learnable triplane tokens, we further adopt Geometry-aware Cross Attention to achieve the image feature injection in a geometry-aware manner. Finally, we utilize the NeRF [69] renderer to generate 3D contents with the generated triplane token.
III-A Geometry-aware Positional Embeddings
@@
In LRM [61] and its variants [70, 65], the triplane tokens are randomly initialized without any priors and directly optimized via back propagation. For convenience, we denote the randomly initialized triplane tokens as $\mathbf{T}_{r}$. In this section, we propose Geometry-aware Positional Embedding (GaPE) to better utilize image prior instead of the randomly initialized triplane tokens.
@@
Geometry Volume Construction. Inspired by the previous multi-view reconstruction method [48], we aim to directly learn a geometry volume from the input images which encodes local geometric information based on the hints of multi-view consistency of the images. To cope with this, we first construct a grid volume whose vertices $\mathbf{X}=\{\mathbf{x}^{j}\}$ are uniformly sampled in the region of interest to be reconstructed, where $\mathbf{x}^{j}$ denotes the $j$-th vertex of the spatial grid volume.
Considering the $i$-th view among the $N$ input views, with its intrinsic $K_{i}\in\mathcal{R}^{3\times 3}$ and extrinsic $E_{i}\in\mathcal{R}^{3\times 4}$, we back project the vertices to the feature map of the $i$-th view $F_{i}$, and derive the corresponding feature grid ${G}_{i}=\{F_{i}(\pi_{i}(\mathbf{x}^{j}))\}$ for the $i$-th view, where $\pi_{i}$ indicates the projection that projecting a 3D vertex onto the $i$-th image plane. The projection $\pi_{i}$ is defined as
@@
where $\mathbf{x}^{j}\in\mathcal{R}^{3\times 1}$ is the 3D coordinate of the grid vertex and $\tilde{\mathbf{x}}^{j}\in\mathcal{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{x}^{j}$. $\mathbf{u}^{j}_{i}\in\mathcal{R}^{2\times 1}$ is the 2D coordinate of 2D image space and $\tilde{\mathbf{u}}^{j}_{i}\in\mathcal{R}^{3\times 1}$ is the homogeneous coordinate of $\mathbf{u}_{i}^{j}$. With the obtained 2D coordinates, we could sample corresponding features from the 2D feature map $F_{i}$. After back-projecting all the grid vertices, we obtain $N$ feature grids $\{G_{i}\}_{i=1,\cdot\cdot\cdot,N}$, which are fused together into one unified feature volume $\mathbf{G}$ with a 3D convolution network.
@@
Figure 3: Pipeline of Geometry-aware Positional Embedding(GaPE) and Geometry-aware Cross Attention(GCA). GaPE (top) fuse the enriched priors of input views in the grid features and randomly initialized triplane tokens to make the triplane tokens geometrically aware. GCA (bottom) use triplane tokens as query and find the spatially corresponding tokens in each feature grid as key and value to incorporates 3D prior and interrelations of the multiple input views.
III-B Geometry-aware Cross Attention
Figure 4: The illustration of Geometry-aware Cross Attention. The vertices of geometry volume are projected back onto the 2D images. The features corresponding to the projected 2D locations are sampled as key and value in GCA.
In addition to the geometry-aware triplane tokens, we introduce a novel cross-attention mechanism, named Geometry-aware Cross Attention (GCA), which incorporates explicit 3D-aware priors for enhanced information fusion between image features and triplane tokens. Specifically, unlike the vanilla attention used in previous LRM works [61, 70, 65] that accesses global features from all views, our method constructs an multi-view geometry-aware attention mechanism. This mechanism ensures that each triplane token correlates with only a limited number of image feature tokens, improving computational efficiency and relevance.
Intuitively, each token in a triplane corresponds to a ray in the 3D space passing through the token, which can be projected into each view (as illustrated in Fig. 4). Consequently, when performing cross-attention between each triplane token and the image features, we only need to correlate the triplane token with the feature tokens intersecting each projected ray. In prior works, dense cross-attention operates on $N\times h\times w$ feature tokens, where $N$ is the number of views, and $h$ and $w$ are the height and width of the feature map. In contrast, our geometry-aware attention operates on only $N\times V$ feature tokens, where $V$ is the size of the geometry volume. This significantly reduces the computational complexity and focuses attention on more relevant features.
@@ III-B Geometry-aware Cross Attention
Specifically, using a triplane token $t$ as query, we gather all vertices of the grid volume that is corresponding to $t$, then project these vertices to every 2D images as illustrated in Figure 4. Features are sampled according to the projected 2D locations and are used as key and value in our GCA module.
To preserve global information interaction among different triplane tokens, we only replace half of the original transformer blocks with GCA.
TABLE I: Comparison of multi-view reconstruction with state-of-the-art image-to-3D counterparts on the Google Scanned Objects (GSO) dataset. $\uparrow$ means the higher the better, and $\downarrow$ indicates the lower the better.
@@
Instant3D Objaverse-LVIS 256 22.18 0.958 0.142 0.0055 0.7012 0.9167
LGM Objaverse 512 21.74 0.967 0.110 0.0034 0.7108 0.9273
Ours (Tiny) Objaverse-LVIS 256 24.35 0.961 0.142 0.0021 0.7327 0.9491
Ours (Base) Objaverse 512 25.23 0.977 0.096 0.0011 0.8725 0.9695
III-C Training Objectives
After obtaining the decoded triplane, we adopt NeRF to render colors $\hat{I}$ and mask images $\hat{m}$. The predicted novel views are supervised via differentiable rendering by other sampled views. We adopt MSE loss, BCE mask loss and LPIPS loss [71]. The training loss can be formulated as
$\begin{gathered}\mathcal{L}_{mse}=\sum\limits_{i=1}^{N}\Big{|}\Big{|}\hat{I}_{i}-I_{i}\Big{|}\Big{|}_{2}\\ \mathcal{L}_{lpips}=\sum\limits_{i=1}^{N}\mathcal{L}_{lpips}(\hat{I}_{i},I_{i})\\ \mathcal{L}_{mask}=\sum\limits_{i=1}^{N}m\log\hat{m}+(1-m)\log(1-\hat{m})\\ \mathcal{L}_{NeRF}=\lambda_{1}\mathcal{L}_{mse}+\lambda_{2}\mathcal{L}_{lpips}+\lambda_{3}\mathcal{L}_{mask}\end{gathered}$ (3)
where $I_{i}$ and $m_{i}$ denote the i-th rendered ground truth images and $\lambda_{i}~{}$(i=1,2,3) are the corresponding weight of MSE, LPIPS and mask loss. The coefficients for different loss components are $\lambda_{1}=1.0$, $\lambda_{2}=0.1$ and $\lambda_{3}=1.0$.